OpenStudy (anonymous):

Which is the radius of a circle whose equation is (x + 1)2 + (y - 1)2 = 25? A 30 B 25 C 5 D 4 Answer and explanation please.

OpenStudy (anonymous):

Equation of a circle is of the form: (x-a)^2 + (y-b)^2 = r^2 Where (a,b) is the center and r is the radius. In this case, 25 is r^2. Therefore, r = 5.

OpenStudy (anonymous):

Thanks for the explanation.